OpenStudy (anonymous):

A rocket starts from rest and moves upward from the surface of the earth. For the first 10s of its motion, the vertical acceleration of the rocket is given by ay=(3.00m/s3)t, where the +y-direction is upward.

OpenStudy (anonymous):

Equations: D=Do + VoT + .5aT^2 V=Vo +aT Height of the rocket will be D= .5*3*10^2 150 meters after 10 seconds. To calculate the speed 200 = .5*3*T^2 200/(.5*3)=T^2 sqrt(200/(.5*3))=T T=11.54 V= 0 +3*11.54 34.641 m/s

OpenStudy (anonymous):

Try that^